The conversion for Celsius to Fahrenheit temperatures is F=9/5C+32. Rewrite the formula so that C is given in terms of F. Then find the Celsius temperature when the Fahrenheit temperature is 81 degrees.
Answer by CharlesG2(834)

The conversion for Celsius to Fahrenheit temperatures is F=9/5C+32. Rewrite the formula so that C is given in terms of F. Then find the Celsius temperature when the Fahrenheit temperature is 81 degrees.
F = 9/5 * C + 32
F - 32 = 9/5 * C
5/9 * (F - 32) = C (this is the formula to solve for C Celsius)
now plug in 81 degrees for F
5/9 * (81 - 32) = C
5/9 * 49 = C
27 2/9 = C
C is approx 27.2 degrees Celsius